OpenStudy (anonymous):

a=1/2h(b1+b2) solve for b2

OpenStudy (anonymous):

\[a=\frac{1}{2}h(b1+b2)\] \[\frac{2a}{h}=b1+b2\] \[b2=\frac{2a}{h}-b1\]